Hello we are doing substitution and i am having trouble doing it with fractions. i will use this for my fraction sign /. please help.
x/2+y/3=7/6
x/4-3y/2=9/4

Answer by stanbon(75887)

Hello we are doing substitution and i am having trouble doing it with fractions. i will use this for my fraction sign /. please help.
x/2+y/3=7/6
---
Get rid of the fractions:
Multiply thru by 6 to get
3x + 2y = 7
-----
x/4-3y/2=9/4
Multiply thru by 4 to get:
x - 6y = 9
--------------------------------
New system:
3x+2y=7
x-6y=9
--------------------
Substitute for "x" and solve for "y":
3(6y+9)+2y = 7
18y + 27 + 2y = 7
20y = -20
y = -1
--------------
Substitute into x = 6y+9 to solve for "x":
x = 6*-1+9
x = -6+9
x = 3
--------------
Cheers,
Stan H.